OpenStudy (anonymous):

Divide the combinations. 60C3 / 15C3

OpenStudy (anonymous):

When I find the two combinations, and then divide them, I get an answer that isn't in the answer options.

OpenStudy (anonymous):

before you start multiplying make sure to cancel your butt off

OpenStudy (anonymous):

I'll solve it here. Let me get the calc.

OpenStudy (anonymous):

60*59*58 3*2*1 4* 59 * 29 -------- * ------ = --------- 3*2*1 15*14*13 7 * 13

OpenStudy (anonymous):

That's the answer I get. Not an answer in the options.

OpenStudy (anonymous):

75.208791208791208791208791208791

OpenStudy (anonymous):

60*59*58/15*14*13

OpenStudy (anonymous):

\[\frac{60\times 59\times 58}{3\times 2}\times \frac{3\times 2}{15\times 14\times 13}\]

OpenStudy (anonymous):

I got the answer. Thank you.

OpenStudy (anonymous):

\[\frac{6844}{91}\]?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yup - what you said is corect